Q1) Evaluate the partial derivative with respect to x of each of the following functions d) f(x, y)2y In(x) e) f(x, y) = 2x-2-tu
Answer #1

C) f(x,y) = 3y/2x2

Partial derivative wrt "X"

f(x,y) = 3y d(1/2x2)/dx

f(x,y) = 3y (-2 x(-2-1)/2)

f(x,y) = -3y/x​​​​​​3​​​

d) f(x,y) = 2yx ln(x)

Partial derivative wrt x

f(x,y) = 2y d(x lnx)

f(x,y) = 2y (x (d(lnx)/dx) + lnx (d(x)/dx )

f(x,y) = 2y ( x(1/x)+ lnx)

f(x,y) = 2y (1+lnx)

e) f(x,y) = (2x+y-yx​​​2​​)/2x

f(x,y) = 1 - (yx/2) + (y/2x)

Partial derivative wrt x

f(x,y) = (d(1)/dx )+ (d(y/2x)) -( d(yx/2)/dx)

f(x,y) = 0 + (-y x​​​​​(-1-1)/2) - (y/2)

f(x,y) = (--y/2x2)+(--y/2)